Given the nodes x0 < x1 < · · · < xn, let V be the vector space of functions that are twice continuously differentiable at each node xi and cubic polynomial on (−∞, x0), (xn, ∞), and each of the intervals (xi , xi+1). Show that any function s(x) ∈ V can be uniquely represented as

s(x) = a0 + a1x + a2x^2 + a3x^3 + "summation" (i=0 to n) ci (x − xi) ^3 +

where (x − xi)^3+ is 0 for x ≤ xi and (x − xi)^3 otherwise. Conclude that this vector space has dimension n + 5.

1
Expert's answer
2021-05-25T15:33:54-0400

Function "s(x)" can be twice differentiated:

"s'(x)=a_1+2a_2x+3a_3x^2+3\displaystyle\sum^n_{i=0}c_i(x-x_i)^2_+"

"s''(x)=2a_2+6a_3x+9\displaystyle\sum^n_{i=0}c_i(x-x_i)_+"

Also, by the theorem of Existence and uniqueness of the polynomial interpolant:

Given the data pairs "\{x_j,f_j\}^n_{j=0}" , where the points "\{x_j\}^n_{j=0}" are distinct, there exists a unique polynomial "p_n(x)" that satisfies the interpolation conditions "p_n(x_i)=f_i,\ i=0,1,2,...,n"

So, any function s(x) ∈ V can be uniquely represented as

"s(x)=a_0+a_1x+a_2x^2+a_3x^3+\displaystyle\sum^n_{i=0}c_i(x-x_i)^3_+"